# EdgeGaussians - 3D Edge Mapping via Gaussian Splatting

### v1

@@ Appendix B Qualitative Results over Replica
Appendix C Limitations / Failure Cases
We now review the limitations of the proposed method.
Fine structures.  In the qualitative results, we observe that the proposed method can not always reconstruct the edges of thin or near planar structures. Fig. 10 shows 3 such examples on the ABC-NEF [91] dataset where the method generates either broken curves or single edges in presence of fine multiple edges. The object on the right in Fig. 10 further has edges inside the object that are not detected in the 2D edge maps, i.e., there is no supervision for those edges, hence it is not possible to map them in 3D.
Large Scale Scenes.  The method has been tested with bounded scenes scaled within a unit cube. Scaling robustly to unbounded scenes is left for future work.
@@ Appendix C Limitations / Failure Cases
(glossaries) Package glossaries Warning: No “printglossary or “printglossaries found. Ω(Remove “makeglossaries if you don’t want any glossaries.) ΩThis document will not have a glossary

### v2

@@ Appendix B Additional Qualitative Results
Appendix C Limitations and Failure Cases
Fine structures and geometric regularization.  As briefly described in the main paper, the method is limited by the noise in the supervisory signal of the 2D edge maps. In many cases the fine structures in such edge maps [66, 57] are not discernible, leading to incorrectly positioned edge points. Geometric regularization applied to noisy edge points can lead the Gaussians’ to form short local curves to satisfy the alignment with their nearest neighbors. Examples of such cases can be seen in Fig. 11.
Clustering and edge fitting.  Further, the clustering algorithm exhibits limitations when applied to larger scenes with complex structures. Fig. 10 shows examples from the Tanks and Temples dataset [36] where the oriented edge points (red), i.e., the 3D Gaussians, better cover the ground-truth 3D edges than the paramtric edges (black). One explanation is that the graph traversal based clustering removes several correct edge components close to the true scene structure while including several incorrect edges. Instead of relying only on local geometric heuristics, defining a prior on which parts of the scene are more likely to hold 3D edge could improve the method’s robustness.
@@ Appendix C Limitations and Failure Cases
Figure 10: Tanks and Temples [36]: Qualitative result showing edges produced by our method on three scenes from the tanks and temples dataset. Supervisory signal (Left), edge points represented as a small line segment centered at the mean of the optimized 3D Gaussians and oriented towards their principal directions (Middle) and the points sampled on the parametric 3D edges estimated (Right). Note that the estimated Gaussians faithfully represent the scene but the clustering and edge fitting process have room for improvement as many correct edges are missed and spurious ones are created in this process.